Budget Performance Report

McAlisters Bottle Company manufactures plastic two-liter bottles for the beverage industry. The cost standards per 100 two-liter bottles are as follows:

Cost Category Standard Cost
per 100 Two-Liter
Bottles
Direct labor $1.52
Direct materials 5.40
Factory overhead 0.40
Total $7.32

At the beginning of May, McAlisters Bottle’s management planned to produce 580,000 bottles. The actual number of bottles produced for May was 630,000 bottles. The actual costs for May of the current year were as follows:

Cost Category Actual Cost for the
Month Ended May 31
Direct labor $9,770
Direct materials 33,200
Factory overhead 2,550
Total $45,520

a. Prepare the May manufacturing standard cost budget (direct labor, direct materials, and factory overhead) for McAlisters Bottle Company, assuming planned production. Enter all amounts as positive numbers.

McAlisters Bottle Company
Manufacturing Cost Budget
For the Month Ended May 31
Standard Cost
at Planned Volume
(580,000 Bottles)
Manufacturing costs:
Direct labor $
Direct materials
Factory overhead
Total $

b. Prepare a budget performance report for manufacturing costs, showing the total cost variances for direct materials, direct labor, and factory overhead for May. Enter favorable variances as negative numbers. Enter unfavorable variances and all other amounts as positive numbers.

McAlisters Bottle Company
Manufacturing Costs - Budget Performance Report
For the Month Ended May 31


Actual Costs
Standard Cost
at Actual Volume
(630,000 bottles)
Cost Variance
(Favorable)
Unfavorable
Manufacturing costs:
Direct labor $ $ $
Direct materials
Factory overhead
Total manufacturing cost $ $ $

c. Which of the following statements is true regarding McAlisters's budget performance report?
The Actual Production Volume was less than planned

The only favorable variance was was related to direct labor

The only unfavorable variance was related to direct materials

Actual Total Costs were less then the amount budgeted
Answer #1
a
McAlisters Bottle Company
Manufacturing Cost Budget
For the Month Ended May 31
Standard Cost
at Planned Volume
(580,000 Bottles)
Manufacturing costs:
Direct labor 8816 =5800*1.52
Direct materials 31320 =5800*5.40
Factory overhead 2320 =5800*0.40
Total 42456
b
McAlisters Bottle Company
Manufacturing Costs - Budget Performance Report
For the Month Ended May 31
Actual Costs Standard Cost Cost Variance
at Actual Volume (Favorable)
(630,000 bottles) Unfavorable
Manufacturing costs:
Direct labor 9770 9576 194
Direct materials 33200 34020 -820
Factory overhead 2550 2520 30
Total manufacturing cost 45520 46116 -596
c
Actual Total Costs were less then the amount budgeted
Workings:
Standard cost:
Manufacturing costs:
Direct labor 9576 =6300*1.52
Direct materials 34020 =6300*5.40
Factory overhead 2520 =6300*0.40